From: NASA HQ/SpaceRef/NASA Watch
Posted: Tuesday, June 27, 2006
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2006) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
The crew took the CHeCS emergency medical operations OBT (on-board training) drill, a 30-min. exercise to refresh the Crew Medical Officer (CMO)'s acuity in applying ACLS (advanced cardio life support) in an emergency. [The training drill, each crewmember for himself, refreshes their memory of the on-orbit stowage and deployment locations, equipment etc. and procedures.]
FE/SO Jeff Williams began his second NASA/JSC renal (kidney) stone session (of three planned), starting his diet log and later setting up the experiment hardware for the 24-hr. void-by-void urine collection starting tomorrow morning and ending on Sunday morning (6/25). [This long-range preventive medicine investigation features daily random ingestion of either potassium citrate or placebo tablets. It is Dr. Peggy Whitson's double-blind research study investigating methods to prevent formation of kidney stones in zero-G. Part of the experiment consists in keeping a metabolic diet log (food and fluid intake), followed by collection of samples several times per day.]
Pavel Vinogradov prepared the fourteenth batch of raw data (broadcast sync commands & message files) from the ongoing testing of the ASN-M satellite navigation system by TsUP, for subsequent downlink via U.S. OCA assets. The 28-day test period came to an end today, and the CDR closed down all data applications, deactivated the LT3 laptop and tore down the data gathering setup. [ASN-M will be required for the arrival of the European ATV (Automated Transfer Vehicle) "Jules Verne" next year. During the extensive checkout from ESA (European Space Agency)/Oberpfaffenhofen, special software on Laptop 3 (running the "Solaris" Operating System) was used for onboard storage of test data and logs. MCC-H had created a special subfolder for ASN data on the OCA file server to facilitate subsequent downlinking of the ASN-M output data, which was performed every other day throughout the 28-day test period for comparison with the uplinked data. The objective of this test was to verify compliance of the ASN-M data with approved requirements and to update ESA's ASN-M model.]
Williams performed maintenance work on the aft IMV (Intermodular Ventilation) fan in the Node.
CDR Vinogradov conducted a visual assessment of free stowage spaces in the Russian segment (RS), recording his observations of available room in the FGB, DC1 Docking Compartment and Service Module (SM) in a log table for subsequent downlink TsUP/Moscow.
At ~5:25am EDT, the CDR also linked up with TsUP specialists to conduct the weekly IMS tagup, discussing open issues concerning identification of equipment and stowage locations for the IMS databases via S-band. [Topics of discussion included open questions from last week's stowage activities, equipment relocations for EVA-16, and amount and location of trash accumulated since closure and undocking of Progress 20, etc.]
After FE Williams configured Lab comm equipment for TV downlink of Progress M-57/22P docking, the crew activated the SSC (Station Support Computer) A31p laptop in the FGB and checked out the functionality of the Russian video system in the SM with the Ku-band equipment in the US segment (USOS). Later the function of the video downlink was also tested and the photo/TV laptop deactivated, leaving cable connections in place. [The A31p is used for routing the video signal from the SM through the FGB via coaxial cable to the SSC OpsLAN (Operations Local Area Network) and from there into the Ku-band system for subsequent conversion from the Russian SECAM format to the American NTSC format on the ground.]
Jeff had another two hours set aside for continuing to collect and prepack hardware slated for return on ULF1.1.
Meanwhile, Pavel similarly gathered and prepacked Russian equipment for the Shuttle, particularly two root modules of the BIO-5 Rasteniya-2 ("Plants-2") experiment's Lada-8 "greenhouse" that was operated earlier this year to study plant growth in micro-G.
The FE performed the regular atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen Sensor) and CDMK (CO2 Monitoring Kit).
Pavel Vladimirovich completed the routine daily maintenance of the SOZh environment control & life support system in the SM, including the ASU toilet facilities systems/replaceables, plus the weekly inspection of the BRPK air/liquid condensate separator apparatus.
Later Pavel also updated/edited the standard IMS (Inventory Management System) "delta file", including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Both crewmembers worked out in their regular 2.5-hr. physical exercise program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (FE, CDR), RED (FE) and VELO bike with bungee cord load trainer (CDR). [Pavel Vinogradov's daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today: Day 1 of the first set).]
Afterwards, Jeff transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
Still showing on Pavel's "time permitting" discretionary task list for today was the search for some missing pieces of a complement of 26 CO2 filter units and six KL-152 TV cameras listed for return to Earth on ULF1.1.
At ~4:05am EDT, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.
At ~3:15pm tonight, the crew will conduct their regular weekly tagup with the Lead Flight Director at MCC-Houston.
MCC-H conducted a ground-commanded leak check of a VES QD (Vacuum Exhaust System Quick Disconnect) in the Lab (LAB1O3). [Other QDs have exhibited some small leaks in the past. A verification of the leak rate of LAB1O3 is desired because this QD location is being considered as an alternate vacuum source for the MSSR-1 materials science rack.]
Yesterday, CMG-3 (Control Moment Gyroscope 3) exhibited offnominal signatures in SMCC (Spin Motor Commanded Current) and vibrations, first appearing coincidental with the SSRMS (Space Station Remote Maneuvering System) checkout. They stabilized and returned to normal values after ~5.5 hours. [Such elevated SMCCs and vibrations have been seen before during Robotics ops, but their long persistence is clearly unusual.]
With the Solar Beta Angle dropping below 60 deg tonight, station attitude will routinely be changed (at ~9:47pm EDT) from earth-fixed LVLH YVV (local vertical local horizontal/y-axis in velocity vector) to sun-oriented XPOP TEA (x-axis perpendicular to orbit plane/torque equilibrium attitude), maneuvered by RS thrusters controlled by the USTO (US Thrusters Only) controller (which obviates handovers between the motion control systems of the two segments).
Progress M-57/22P Launch Status (picture below): The Soyuz-U/22P system is undergoing L-1 launch preparations at Baikonur, Kazakhstan. Liftoff is scheduled for tomorrow, at 11:08am EDT (9:08pm at Baikonur). [22P carries 2400 kg of various cargoes, including food, water, fuel, clothes for the E14 crew, and equipment. Pavel and Jeff will also get fresh vegetables, fruit and parcels from their families. In addition, 22P delivers gifts for E13 CDR Vinogradov who will be celebrating his 53rd birthday on August 31. U.S. hardware requested to fly on 22P includes 40 U.S. food containers, clothing items, 2 IMAKs (ISS Medical Accessory Kits), 2 TVIS harnesses, 2 advanced EMU batteries, 2 LCVGs (Liquid Cooling & Ventilation Garments), 6 DST (Dual Sorbent Tube) assemblies and other water, surface & air sampling gear, new renal (kidney) stone experiment supplies, etc.]
Today's CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Mt. Kilimanjaro, Kenya (looking right for this ice-covered peak. The ice caps are projected to melt completely within about ten years), Patagonian Glaciers (the southern end of this site, especially on the dry eastern side, should have been clear. Images of smaller glacier tongues at lower altitudes were requested), and South Amazonian Fans, Brazil (with unusually clear skies over the Amazon basin, detailed imagery is requested to study a new class of features more closely. Handheld imagery from STS94 indicated the unexpected existence of large fans [partial cones of river sediment, >100 km radius] in the Amazon basin, hundreds of km from the Andean mountain front [where numerous such features exist]. This discovery allowed scientists to propose the existence of large fans along the flanks of the Amazon basin, where many have since been located. Ancient river patterns can be seen as a tracery of wavy light colored lines in the rainforest. These are sandy river courses which yield such poor soils that tree growth is significantly reduced, and river patterns can be seen from low Earth orbit).
Note: During the next week or two, ISS daylight "awake" passes will be located primarily in the Southern Hemisphere where lighting has just passed its seasonal minimum (solstice, 6/21) and winter weather will further reduce the crew's viewing opportunities. Station tracks for now are also paralleling the terminator and sun angles will remain low. There will probably be a number of days where no suitable CEO target will be available for daily uplink.
To date, more than 198,000 of CEO images have been taken in the first five years of the ISS, almost one third of the total number of images taken from orbit by astronauts.
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 13 crew visit:
Expedition 13 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit (as of this morning, 7:14am EDT [=3D epoch]):
Significant Events Ahead (all dates Eastern and subject to change):
// end //